OpenStudy (anonymous):

17. If you are given the length of sides a and b and the m

OpenStudy (anonymous):

17. Law of sines \[\frac{ a }{ \sin A }=\frac{ b }{ \sin B}=\frac{ c }{ \sin C }\] so i would say B. No. You would need to know the m<B or the length of side c to use the Law of Sines. 18. again \[\frac{ a }{ \sin A }=\frac{ b }{ \sin B}=\frac{ c }{ \sin C }\] \[\frac{ a }{ \sin A }=\frac{ c }{ \sin C }\] \[\frac{ 10 }{ \sin 33^{o} }=\frac{ c }{ \sin C }\] you must know the value of the side c to find it so D. Not enough information given.

OpenStudy (anonymous):

thanks so much